OpenStudy (anonymous):

Will Medal! TRUE or FALSE? De Moivre's theorem is used to raise a complex number to ANY integer power.

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

what do you know about DE Moivre's theorem

OpenStudy (anonymous):

I find it difficult to understand. haha But just the formula I somewhat understand.

OpenStudy (anonymous):

DeMoivre’s Theorem is used to raise complex numbers to integer powers. DeMoivre’s Theorem: If z = a + bi is any complex number with trig form r cis q and n is any positive integer then the n th power of z is given by z n = r n cis n q = r n ( cos n q + i sin n q ). Example: Find [ 2(cos 15° + i sin 15° )] 6 . Write the result in standard form. [ 2(cos 15° + i sin 15° )] 6 = (2 cis 15° ) 6 = 2 6 cis 6( 15 ° ) = 64 cis 90° = 64 ( cos 90 + i sin 90) = 64 ( 0 + 1 * i ) = 64 i .

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

I notice that it says "positive integer". Does that mean that it will be false because it cannot be negative?